how many liters of a 30% alcohol solution must be mixed with 80 liters of a 80% solution to get a 50 % solution?
Answer by mukhopadhyay(490)

Assume x liters of a 30% alcohol solution is to be mixed with 80 liters of a 80% solution to produce (x+80) liters of a 50% solution.
=> x*(30%) + 80(80%) = (x+80)(50%)
=> .3x + 64 = .5x + 40
=> .2x = 24
=> x = 120
Answer: 120 liters